2. For the gas-phase reaction 2 NH3 <=> N2 + 3 H2, a closed system initially contained 1.00 mol of NH3, 4.50 mol of N2, and 4physical chemistry, please show all work and explanations clearly
Answer #1

The Greek symbol \xi denotes the extent of reaction and is denoted by

Nin,i – neq, =

Here, neq,i demotes equilibrium concentration of i-th reactant and nin,i denotes the initial concentration. {\nu_i} here is the stoichiometric coefficient.

a) Maximum value of \xi will be achieved when all of NH3 will be converted to the reactant and neq for NH3 becomes 0. nin for NH3 will be 1 mole and {\nu_i} i.e. the stoichiometric coefficient is 2. Thus, we have,

1-0 Smar = =0.5

For minimum value of \xi , the reaction should happen in reverse and maximum possible amount of NH3 be formed. For the reverse reaction, 1 mole of N2 will need 3 moles of H2 and 4.50 moles of N2 will need 3*4.50 = 13.50 moles of H2 but there are only 4.20 moles. Hence, for the reverse reaction, H2 will be the limiting reagent and it will go down to 0.

3 moles H2 will form 2 moles NH3. Thus, 4.20 moles of H2 will form 4.20*2/3=2.80 moles NH3. But 1.00 mole NH3 was already there in the system.

Thus, for minimum \xi , we have, neq for NH3 will be 2.80+1.0 =3.80.

1-3.8 Smin = = -1.4

Thus, finally,

-1.40 < <0.5

b)

Here, we need to calculate maximum and minimum moles of each component.

Maximum extent of forward reaction will give minimum moles of NH3 and maximum moles for H2 and N2

For NH3,

піn,i — пеgi 1.00 – пеа

Putting {= Smar = 0.50 , we will have,

1.00 – neg 0.50 = neg=0.00

Putting <= min = -1.40 , we will have,

1.00 - ne -1.40 = = neg = 3.80

Thus, minimum possible value of nNH3 is 0.00 moles and maximum possible value is 3.80 moles.

For N2.

Here, nin = 4.50 moles and vi =1

Here, because Nitrogen is on product side, we will use the add a negative sign.

* = Mingi – Negi – 4.50 – neq

Putting {= Smar = 0.50 , we will have,

4.50 - ne 0.50 = - 1 neg= 5.00

Putting <= min = -1.40 , we will have,

-1.40 = _ 4.50 – neg neg= 3.10

Thus, minimum possible value of nN2 is 3.10 moles and maximum possible value is 5.00 moles.

For H2

Here, nin = 4.20 moles and vi =3

Here, because Hydrogen is on product side, we will use the add a negative sign.

піn,i — пеgi 4.20 – пеа S = --

Putting {= Smar = 0.50 , we will have,

4.20 – ne 0.50 = = neg= 5.70

Putting <= min = -1.40 , we will have,

4.20 - neg -1.40 = - = neg = 0.00

Thus, minimum possible value of nH2 is 0.00 moles and maximum possible value is 5.70 moles.